Could someone please show me step-by-step how to solve this? Thank You for your help.
Solve the system.
4x^2-4y^2=20
3x^2+2y^2=35

Answer by stanbon(75887)

Solve the system.
4x^2-4y^2=20
3x^2+2y^2=35
--------------------
Multiply the 2nd equation by thru by 2 to get:
6x^2+4y^2=50
---------
Add that to the 1st equation to get:
10x^2 = 70
x = +/-sqrt7
-------------
Substitute into 4x^2-4y^2=20 to solve for "y":
4*7 -4y^2 = 20
4y^2 = 8
y = +/-sqrt2
===============
Cheers,
Stan H.